INFRARED PHOTODISSOCIATION SPECTROSCOPY OF CATIONIC TRANSITION METAL-NITROGEN COMPLEXES

\underline{E.D. PILLAI} and M. A. DUNCAN, University of Georgia, Department of Chemistry, Athens, GA, 30602-2556.

Transition metal-nitrogen complexes of the form TM^{+}(N_{2})_{n} (TM=V, Fe, Nb) are produced in a laser vaporization pulsed nozzle cluster source, size selected with a time-of-flight mass spectrometer and excited in the infrared with an infrared OPO/OPA. Infrared photodissociation occurs by elimination of whole N_{2} molecules and the photofragement yield versus wavelength produces IR resonance-enhanced photodissociation (REPD) spectra for the clusters. Vibrational bands seen in the 2100 - 2400 cm^{−1} region correspond to the N-N stretch but are shifted from those of free N_{2}. Density functional theory calculations on the bond energeries, geometries and frequency shifts for the TM^{+}(N_{2})_{n} are also discussed in regard to periodic trends.